OpenStudy (anonymous):

Fill 'Er gas station stores its various types pf gasoline in understand cylindrical tanks. Each tank has a height of 31 feet and hold 1,337 cubic feet of gasoline when it is full. Find the radius of one tank. Show your work and round answers to hundredths as needed.

OpenStudy (anonymous):

see.. volume of cylinder has formula pi*r*r*h

OpenStudy (anonymous):

Well, a cylinder has circular base and a height h. So it's volume is \[V = \pi r^2 h\]Now we know from the description that h =31ft and we know that V=1337ft. We isolate r in the equation and solve for it.\[r^2 = \frac{V}{\pi h}\]just replace the values and when you get a number, take the square root. And you get the result.

OpenStudy (anonymous):

^ nice one there :P lol

OpenStudy (anonymous):

so would the answer be 114.86?

OpenStudy (anonymous):

Replace it like this,\[r^2=\frac{1337}{31\pi}\]When you calculate that, take the square root. Try again.

OpenStudy (anonymous):

1.77?

OpenStudy (anonymous):

No, its \[r=\sqrt{\frac{1337}{31 \pi}} \approx 3.705\]Try your calculation again. You were close!

OpenStudy (anonymous):

1.92

OpenStudy (anonymous):

No, its 3.705, check my work. And do exactly as I've done to get that value.

OpenStudy (anonymous):

so all i do is square root 1337/ 31 pi. Correct?

OpenStudy (anonymous):

When you do this in a claculator, be very careful. First you have to multiply 31*pi = 97.389, then you take that and use it in 1337/97.389 = 13.7248, and then you take the square root.

OpenStudy (anonymous):

so the final answer is 3.70

OpenStudy (anonymous):

Yeah. Good job!

OpenStudy (anonymous):

do you mind helping with the @nd part of the question?

OpenStudy (anonymous):

How tall would the tank in question 1 have to be in order to triple the volume? Show your work and round answers to hundredths if needed

OpenStudy (anonymous):

Ok, now you have your wolume 1337, triple of that would be 3*1337=4011. Repeat the same procedure, but now h is unknown. So we do: \[h=\frac{4011}{3.70^2 \pi} \approx 93.26\]

OpenStudy (anonymous):

3.70^2*pi=43.01

OpenStudy (anonymous):

yes, now put that under 4011, and divide.

OpenStudy (anonymous):

i got 93.26

OpenStudy (anonymous):

Yes, correct!

OpenStudy (anonymous):

Thanks alot!

OpenStudy (anonymous):

You're welcome katie!